OpenStudy (moonlitfate):

An object is projected upward from ground level with an initial velocity of 660 feet per second. (In this exercise, the goal is to analyze the motion of the object during its upward flight.) (a) If air resistance is neglected, find the velocity of the object as a function of time. Use a graphing utility to graph this function. (b) Use the result of part (a) to find the position function. Determine the maximum height attained by the object.

OpenStudy (aum):

Calculus?

OpenStudy (aum):

\[ v = \frac{dy}{dt} = u - gt \\ y = \int (u - gt)dt \]u = 660 ft / sec g = 32 ft / sec^2

OpenStudy (moonlitfate):

(c) If the air resistance is proportional to the square of the velocity, you obtain the equation \[\int\limits \frac{ dv }{ 32+kv^2 } = -\int\limits dt\] where −32 feet per second is the acceleration due to gravity and k is a constant. Find the velocity as a function of time by solving the equation v(t)= _______________________ (d) Use a graphing utility to graph the velocity function v(t) in part (c) for k = 0.003. Use the graph to approximate the time t0 at which the object reaches its maximum height. (Round your answer to two decimal places.) \[t _{0} = ______ s\] (e) Use the integration capabilities of a graphing utility to approximate the integral \[\int\limits_{{0}}^{t_0}v(t) dt\] where v(t) and \[t_{0}\] are those found in part (d). This is the approximation of the maximum height of the object. (Round your answer to two decimal places.) h = ______ ft

OpenStudy (moonlitfate):

Calculus II, not I. ._. @aum

OpenStudy (moonlitfate):

I know for the (a) that v(t) = -32t+660

OpenStudy (moonlitfate):

And that for (b) that the position function is \[-16t^2+660t\]

OpenStudy (aum):

a) \(\large v = u - gt \) \(\large v = 660 - 32t\) 

OpenStudy (moonlitfate):

I'm stuck on how to find maximum height... ._.

OpenStudy (moonlitfate):

Well maximum height of the object...

OpenStudy (aum):

yes, b) \(\large y = -16t^2+660t \) This is a parabola and we need to find the y-coordinate of the vertex. The x-coordinate of the vertex is given by -b/(2a) (for ax^2 + bx + c) -b/2a = -660 / (2 * -16) = ? seconds. put that t in \(\large y = -16t^2+660t \) and find \(\large y_{max}\).

OpenStudy (moonlitfate):

Doing that now.

OpenStudy (moonlitfate):

6806.25= y max

OpenStudy (aum):

Yes, maximum height = 6806.25 feet. Max. height can also be found this way: When max height is reached, velocity will be zero because if the velocity is not zero the object will keep moving upwards. Since it has stopped and about to return, v = 0 at max height. From a) v = 660 - 32t 0 = 660 - 32t t = 20.625 sec. ymax = -16(20.625)^2 + 660(20.625) = 6806.25 feet.

OpenStudy (moonlitfate):

Oh, wow! Got it. :) That makes that part easier.

OpenStudy (aum):

For c) make a tan substitution. Or use the standard integral for \(\large \int \frac{dx}{a^2+x^2}\)

OpenStudy (moonlitfate):

The substution and integration I'm having trouble with. Rusty skills, but I'm trying to work on them.

OpenStudy (moonlitfate):

But for the left side of the equation, the tan substitution can be applied right? I'm trying to understand. :)

OpenStudy (aum):

\[ \int\limits \frac{ dv }{ 32+kv^2 } = -\int\limits dt \\ \frac 1k \int\limits \frac{ dv }{ 32/k+v^2 } = -\int\limits dt \\ \frac 1k \int\limits \frac{ dv }{ \left(\sqrt{32/k}\right )^2+v^2 } = -\int\limits dt \\ v = \left(\sqrt{32/k}\right )\tan(t) \] Or you can look up standard integration table for \[ \int\limits \frac{ dx }{ a^2+x^2 }\]

OpenStudy (aum):

\[\int\limits \frac{ dx }{ a^2+x^2 } = \frac 1a \tan^{-1}\frac xa + C\]

OpenStudy (aum):

gtg.

OpenStudy (moonlitfate):

Okay, thank you for the help though. :)